\begin{thebibliography}{10}
	
	\bibitem{Guth:1980zm}
	Guth A.~H 1981 {The Inflationary Universe: A Possible Solution to the Horizon
		and Flatness Problems} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.23.347}{\textbf {23} 347}
	
	\bibitem{Linde:1981mu}
	Linde A.~D 1982 {A New Inflationary Universe Scenario: A Possible Solution of
		the Horizon, Flatness, Homogeneity, Isotropy and Primordial Monopole
		Problems} {\it Phys. Lett. B}
	\newblock  \href{https://doi.org/10.1016/0370-2693(82)91219-9}{\textbf {108}
		389}
	
	\bibitem{Albrecht:1982wi}
	Albrecht A and Steinhardt P.~J 1982 {Cosmology for Grand Unified Theories with
		Radiatively Induced Symmetry Breaking} {\it Phys. Rev. Lett.}
	\newblock  \href{https://doi.org/10.1103/PhysRevLett.48.1220}{\textbf {48}
		1220}
	
	\bibitem{Starobinsky:1980te}
	Starobinsky A.~A 1980 {A New Type of Isotropic Cosmological Models Without
		Singularity} {\it Phys. Lett. B}
	\newblock  \href{https://doi.org/10.1016/0370-2693(80)90670-X}{\textbf {91} 99}
	
	\bibitem{Guth:1982ec}
	Guth A.~H and Pi S.~Y 1982 {Fluctuations in the New Inflationary Universe} {\it
		Phys. Rev. Lett.}
	\newblock  \href{https://doi.org/10.1103/PhysRevLett.49.1110}{\textbf {49}
		1110}
	
	\bibitem{Bezrukov:2007ep}
	Bezrukov F.~L and Shaposhnikov M 2008 {The Standard Model Higgs boson as the
		inflaton} {\it Phys. Lett. B}
	\newblock  \href{https://doi.org/10.1016/j.physletb.2007.11.072}{\textbf {659}
		703}
	
	\bibitem{Germani:2010gm}
	Germani C and Kehagias A 2010 {New Model of Inflation with Non-minimal
		Derivative Coupling of Standard Model Higgs Boson to Gravity} {\it Phys. Rev.
		Lett.}
	\newblock  \href{https://doi.org/10.1103/PhysRevLett.105.011302}{\textbf {105}
		011302}
	
	\bibitem{Yi:2018gse}
	Yi Z, Gong Y and Sabir M 2018 {Inflation with Gauss-Bonnet coupling} {\it Phys.
		Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.98.083521}{\textbf {98}
		083521}
	
	\bibitem{Germani:2014hqa}
	Germani C, Watanabe Y and Wintergerst N 2014 {Self-unitarization of New Higgs
		Inflation and compatibility with Planck and BICEP2 data} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2014/12/009}{\textbf {12}
		009}
	
	\bibitem{Yang:2015pga}
	Yang N, Fei Q, Gao Q and Gong Y 2016 {Inflationary models with non-minimally
		derivative coupling} {\it Class. Quant. Grav.}
	\newblock  \href{https://doi.org/10.1088/0264-9381/33/20/205001}{\textbf {33}
		205001}
	
	\bibitem{Germani:2010ux}
	Germani C and Kehagias A 2010 {Cosmological Perturbations in the New Higgs
		Inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2010/05/019}{\textbf {05}
		019}
	
	\bibitem{Minamitsuji:2013ura}
	Minamitsuji M 2014 {Solutions in the scalar-tensor theory with nonminimal
		derivative coupling} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.89.064017}{\textbf {89}
		064017}
	
	\bibitem{Sadjadi:2012zp}
	Sadjadi H.~M and Goodarzi P 2013 {Reheating in nonminimal derivative coupling
		model} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2013/02/038}{\textbf {02}
		038}
	
	\bibitem{Zhu:2015lry}
	Zhu Y and Gong Y 2016 {PPN parameters in gravitational theory with nonminimally
		derivative coupling} {\it Int. J. Mod. Phys. D}
	\newblock  \href{https://doi.org/10.1142/S0218271817500055}{\textbf {26}
		1750005}
	
	\bibitem{Yang:2015zgh}
	Yang N, Gao Q and Gong Y 2015 {Inflation with non-minimally derivative
		coupling} {\it Int. J. Mod. Phys. A}
	\newblock  \href{https://doi.org/10.1142/S0217751X15450049}{\textbf {30}
		1545004}
	
	\bibitem{Gong:2017kim}
	Gong Y, Papantonopoulos E and Yi Z 2018 {Constraints on
		scalar\textendash{}tensor theory of gravity by the recent observational
		results on gravitational waves} {\it Eur. Phys. J. C}
	\newblock  \href{https://doi.org/10.1140/epjc/s10052-018-6227-9}{\textbf {78}
		738}
	
	\bibitem{Fu:2019ttf}
	Fu C, Wu P and Yu H 2019 {Primordial Black Holes from Inflation with Nonminimal
		Derivative Coupling} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.100.063532}{\textbf {100}
		063532}
	
	\bibitem{Granda:2019wip}
	Granda L.~N, Jimenez D.~F and Cardona W 2020 {Higgs inflation with non-minimal
		derivative coupling to gravity} {\it Astropart. Phys.}
	\newblock  \href{https://doi.org/10.1016/j.astropartphys.2020.102459}{\textbf
		{121} 102459}
	
	\bibitem{Oliveros:2019xef}
	Oliveros A and Noriega H.~E 2019 {Constant-roll inflation driven by a scalar
		field with nonminimal derivative coupling} {\it Int. J. Mod. Phys. D}
	\newblock  \href{https://doi.org/10.1142/S0218271819501591}{\textbf {28}
		1950159}
	
	\bibitem{Gialamas:2020vto}
	Gialamas I.~D, Karam A, Lykkas A and Pappas T.~D 2020 {Palatini-Higgs inflation
		with nonminimal derivative coupling} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.102.063522}{\textbf {102}
		063522}
	
	\bibitem{Heydari:2021gea}
	Heydari S and Karami K 2022 {Primordial black holes in nonminimal derivative
		coupling inflation with quartic potential and reheating consideration} {\it
		Eur. Phys. J. C}
	\newblock  \href{https://doi.org/10.1140/epjc/s10052-022-10036-2}{\textbf {82}
		83}
	
	\bibitem{Goodarzi:2022wli}
	Goodarzi P 2022 {Anisotropic inflation in non-minimal kinetic coupling model}
	{\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2022/11/052}{\textbf {11}
		052}
	
	\bibitem{Oikonomou:2024tjf}
	Oikonomou V.~K 2024 {Non-minimal derivative coupling theories compatible with
		GW170817} {\it Nucl. Phys. B}
	\newblock  \href{https://doi.org/10.1016/j.nuclphysb.2024.116467}{\textbf
		{1000} 116467}
	
	\bibitem{Gialamas:2024jeb}
	Gialamas I.~D, Katsoulas T and Tamvakis K 2024 {Inflation and reheating in
		quadratic metric-affine gravity with derivative couplings}
	\newblock  arXiv: \href{https://arXiv.org/abs/2403.08530}{2403.08530}
	
	\bibitem{Hell:2024xbv}
	Hell A 2024 {Unveiling the inconsistency of the Proca theory with non-minimal
		coupling to gravity}
	\newblock  arXiv: \href{https://arXiv.org/abs/2403.18673}{2403.18673}
	
	\bibitem{Martin:2012pe}
	Martin J, Motohashi H and Suyama T 2013 {Ultra Slow-Roll Inflation and the
		non-Gaussianity Consistency Relation} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.87.023514}{\textbf {87}
		023514}
	
	\bibitem{Motohashi:2014ppa}
	Motohashi H, Starobinsky A.~A and Yokoyama J 2015 {Inflation with a constant
		rate of roll} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2015/09/018}{\textbf {09}
		018}
	
	\bibitem{Yi:2017mxs}
	Yi Z and Gong Y 2018 {On the constant-roll inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/03/052}{\textbf {03}
		052}
	
	\bibitem{Germani:2017bcs}
	Germani C and Prokopec T 2017 {On primordial black holes from an inflection
		point} {\it Phys. Dark Univ.}
	\newblock  \href{https://doi.org/10.1016/j.dark.2017.09.001}{\textbf {18} 6}
	
	\bibitem{Gong:2017qlj}
	Di H and Gong Y 2018 {Primordial black holes and second order gravitational
		waves from ultra-slow-roll inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/07/007}{\textbf {07}
		007}
	
	\bibitem{Ezquiaga:2017fvi}
	Ezquiaga J.~M, Garcia-Bellido J and Ruiz~Morales E 2018 {Primordial Black Hole
		production in Critical Higgs Inflation} {\it Phys. Lett. B}
	\newblock  \href{https://doi.org/10.1016/j.physletb.2017.11.039}{\textbf {776}
		345}
	
	\bibitem{Ballesteros:2017fsr}
	Ballesteros G and Taoso M 2018 {Primordial black hole dark matter from single
		field inflation} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.97.023501}{\textbf {97}
		023501}
	
	\bibitem{Drees:2019xpp}
	Drees M and Xu Y 2021 {Overshooting, Critical Higgs Inflation and Second Order
		Gravitational Wave Signatures} {\it Eur. Phys. J. C}
	\newblock  \href{https://doi.org/10.1140/epjc/s10052-021-08976-2}{\textbf {81}
		182}
	
	\bibitem{Lin:2020goi}
	Lin J {\it et~al} 2020 {Primordial black holes and secondary gravitational
		waves from $k$ and $G$ inflation} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.101.103515}{\textbf {101}
		103515}
	
	\bibitem{Lin:2021vwc}
	Lin J {\it et~al} 2023 {Primordial black holes and scalar induced gravitational
		waves from Higgs inflation with noncanonical kinetic term} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.107.043517}{\textbf {107}
		043517}
	
	\bibitem{Yi:2020cut}
	Yi Z, Gao Q, Gong Y and Zhu Z.-h 2021 {Primordial black holes and
		scalar-induced secondary gravitational waves from inflationary models with a
		noncanonical kinetic term} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.103.063534}{\textbf {103}
		063534}
	
	\bibitem{Yi:2020kmq}
	Yi Z, Gong Y, Wang B and Zhu Z.-h 2021 {Primordial black holes and secondary
		gravitational waves from the Higgs field} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.103.063535}{\textbf {103}
		063535}
	
	\bibitem{Yi:2021lxc}
	Yi Z and Zhu Z.-H 2022 {NANOGrav signal and LIGO-Virgo primordial black holes
		from the Higgs field} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2022/05/046}{\textbf {05}
		046}
	
	\bibitem{Yi:2022anu}
	Yi Z 2023 {Primordial black holes and scalar-induced gravitational waves from
		the generalized Brans-Dicke theory} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2023/03/048}{\textbf {03}
		048}
	
	\bibitem{Zhang:2021rqs}
	Zhang F 2022 {Primordial black holes and scalar induced gravitational waves
		from the E model with a Gauss-Bonnet term} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.105.063539}{\textbf {105}
		063539}
	
	\bibitem{Wang:2024euw}
	Wang Z, Gao S, Gong Y and Wang Y 2024 {Primordial black holes and
		scalar-induced gravitational waves from the polynomial attractor model} {\it
		Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.109.103532}{\textbf {109}
		103532}
	
	\bibitem{Gao:2019sbz}
	Gao Q, Gong Y and Yi Z 2019 {On the constant-roll inflation with large and
		small $\eta_H$} {\it Universe}
	\newblock  \href{https://doi.org/10.3390/universe5110215}{\textbf {5} 215}
	
	\bibitem{Tzirakis:2007bf}
	Tzirakis K and Kinney W.~H 2007 {Inflation over the hill} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.75.123510}{\textbf {75}
		123510}
	
	\bibitem{Morse:2018kda}
	Morse M. J.~P and Kinney W.~H 2018 {Large-$\eta$ constant-roll inflation is
		never an attractor} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.97.123519}{\textbf {97}
		123519}
	
	\bibitem{Motohashi:2017vdc}
	Motohashi H and Starobinsky A.~A 2017 {$f(R)$ constant-roll inflation} {\it
		Eur. Phys. J. C}
	\newblock  \href{https://doi.org/10.1140/epjc/s10052-017-5109-x}{\textbf {77}
		538}
	
	\bibitem{Oikonomou:2017bjx}
	Oikonomou V.~K 2017 {Reheating in Constant-roll $F(R)$ Gravity} {\it Mod. Phys.
		Lett. A}
	\newblock  \href{https://doi.org/10.1142/S0217732317501723}{\textbf {32}
		1750172}
	
	\bibitem{Odintsov:2017qpp}
	Odintsov S.~D and Oikonomou V.~K 2017 {Inflation with a Smooth Constant-Roll to
		Constant-Roll Era Transition} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.96.024029}{\textbf {96}
		024029}
	
	\bibitem{Nojiri:2017qvx}
	Nojiri S, Odintsov S.~D and Oikonomou V.~K 2017 {Constant-roll Inflation in
		$F(R)$ Gravity} {\it Class. Quant. Grav.}
	\newblock  \href{https://doi.org/10.1088/1361-6382/aa92a4}{\textbf {34} 245012}
	
	\bibitem{Dimopoulos:2017ged}
	Dimopoulos K 2017 {Ultra slow-roll inflation demystified} {\it Phys. Lett. B}
	\newblock  \href{https://doi.org/10.1016/j.physletb.2017.10.066}{\textbf {775}
		262}
	
	\bibitem{Ito:2017bnn}
	Ito A and Soda J 2018 {Anisotropic Constant-roll Inflation} {\it Eur. Phys. J.
		C}
	\newblock  \href{https://doi.org/10.1140/epjc/s10052-018-5534-5}{\textbf {78}
		55}
	
	\bibitem{Karam:2017rpw}
	Karam A, Marzola L, Pappas T, Racioppi A and Tamvakis K 2018 {Constant-Roll
		(Quasi-)Linear Inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/05/011}{\textbf {05}
		011}
	
	\bibitem{Fei:2017fub}
	Fei Q, Gong Y, Lin J and Yi Z 2017 {The reconstruction of tachyon inflationary
		potentials} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2017/08/018}{\textbf {08}
		018}
	
	\bibitem{Cicciarella:2017nls}
	Cicciarella F, Mabillard J and Pieroni M 2018 {New perspectives on
		constant-roll inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/01/024}{\textbf {01}
		024}
	
	\bibitem{Anguelova:2017djf}
	Anguelova L, Suranyi P and Wijewardhana L. C.~R 2018 {Systematics of Constant
		Roll Inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/02/004}{\textbf {02}
		004}
	
	\bibitem{Gao:2018tdb}
	Gao Q, Gong Y and Fei Q 2018 {Constant-roll tachyon inflation and observational
		constraints} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/05/005}{\textbf {05}
		005}
	
	\bibitem{Gao:2018cpp}
	Gao Q 2018 {The observational constraint on constant-roll inflation} {\it Sci.
		China Phys. Mech. Astron.}
	\newblock  \href{https://doi.org/10.1007/s11433-018-9197-2}{\textbf {61}
		070411}
	
	\bibitem{Mohammadi:2018wfk}
	Mohammadi A, Saaidi K and Sheikhahmadi H 2019 {Constant-roll approach to
		non-canonical inflation} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.100.083520}{\textbf {100}
		083520}
	
	\bibitem{Pattison:2018bct}
	Pattison C, Vennin V, Assadullahi H and Wands D 2018 {The attractive behaviour
		of ultra-slow-roll inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2018/08/048}{\textbf {08}
		048}
	
	\bibitem{Lin:2019fcz}
	Lin W.-C, Morse M. J.~P and Kinney W.~H 2019 {Dynamical Analysis of Attractor
		Behavior in Constant Roll Inflation} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2019/09/063}{\textbf {09}
		063}
	
	\bibitem{Fei:2020jab}
	Fei Q, Yi Z and Yang Y 2020 {The Reconstruction of Non-Minimal Derivative
		Coupling Inflationary Potentials} {\it Universe}
	\newblock  \href{https://doi.org/10.3390/universe6110213}{\textbf {6} 213}
	
	\bibitem{Antoniadis:2020dfq}
	Antoniadis I, Lykkas A and Tamvakis K 2020 {Constant-roll in the Palatini-$R^2$
		models} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2020/04/033}{\textbf {04}
		033}
	
	\bibitem{Guerrero:2020lng}
	Guerrero M, Rubiera-Garcia D and Saez-Chillon~Gomez D 2020 {Constant roll
		inflation in multifield models} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.102.123528}{\textbf {102}
		123528}
	
	\bibitem{Shokri:2021iqp}
	Shokri M, Setare M.~R, Capozziello S and Sadeghi J 2022 {Constant-roll f(R)
		inflation compared with cosmic microwave background anisotropies and
		swampland criteria} {\it Eur. Phys. J. Plus}
	\newblock  \href{https://doi.org/10.1140/epjp/s13360-022-02874-6}{\textbf {137}
		639}
	
	\bibitem{Fei:2022gfp}
	Fei Q, Ahmed W and Wang Z.-L 2022 {On the Constant-Roll Tachyon Inflation with
		Large and Small $\eta_{H}$} {\it Symmetry}
	\newblock  \href{https://doi.org/10.3390/sym14122670}{\textbf {14} 2670}
	
	\bibitem{Stojanovic:2023dni}
	Stojanovic M, Bilic N, Dimitrijevic D.~D, Djordjevic G.~S and Milosevic M 2023
	{Constant-roll inflation with tachyon field in the holographic braneworld}
	\newblock  arXiv: \href{https://arXiv.org/abs/2401.00352}{2401.00352}
	
	\bibitem{Huang:2024xqk}
	Huang Q, Huang H, Xu B and Zhang K 2024 {Constant-roll inflation and primordial
		black holes with Barrow holographic dark energy}
	\newblock  arXiv: \href{https://arXiv.org/abs/2401.15451}{2401.15451}
	
	\bibitem{Germani:2011ua}
	Germani C and Watanabe Y 2011 {UV-protected (Natural) Inflation: Primordial
		Fluctuations and non-Gaussian Features} {\it JCAP}
	\newblock  \href{https://doi.org/10.1088/1475-7516/2011/07/031}{\textbf {07}
		031}
	
	\bibitem{Kinney:2005vj}
	Kinney W.~H 2005 {Horizon crossing and inflation with large eta} {\it Phys.
		Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.72.023515}{\textbf {72}
		023515}
	
	\bibitem{Stewart:1993bc}
	Stewart E.~D and Lyth D.~H 1993 {A More accurate analytic calculation of the
		spectrum of cosmological perturbations produced during inflation} {\it Phys.
		Lett. B}
	\newblock  \href{https://doi.org/10.1016/0370-2693(93)90379-V}{\textbf {302}
		171}
	
	\bibitem{Tsujikawa:2012mk}
	Tsujikawa S 2012 {Observational tests of inflation with a field derivative
		coupling to gravity} {\it Phys. Rev. D}
	\newblock  \href{https://doi.org/10.1103/PhysRevD.85.083518}{\textbf {85}
		083518}
	
\end{thebibliography}

\documentclass[aps,prd,showpacs,nofootinbib,reprint]{revtex4-2}

\usepackage{booktabs}
\usepackage{stmaryrd}
\usepackage{mathrsfs}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{comment}
\usepackage{color}
\usepackage{hyperref}
\hypersetup{
  colorlinks=true,
  linkcolor=red,
  citecolor=blue,
}

\begin{document}
\title{Constant-roll inflation with non-minimally derivative coupling}

\author{Jie Liu}
\email{ liujie@hainnu.edu.cn} 
\affiliation{School of Physics and Electronic Engineering, Hainan Normal University, Haikou 571158, China}
\author{Yungui Gong}
\email{gongyungui@nbu.edu.cn} 
\affiliation{Department of Physics, School of Physical Science and Technology, Ningbo University, Ningbo, Zhejiang 315211, China}
\affiliation{School of Physics, Huazhong University of Science and Technology, Wuhan, Hubei
430074, China}
\author{Zhu Yi}
\email{Corresponding author. yz@bnu.edu.cn}
\affiliation{Faculty of Arts and Sciences, Beijing Normal University, Zhuhai 519087, China}
\affiliation{Advanced Institute of Natural Sciences, Beijing Normal University, Zhuhai 519087, China}
\begin{abstract}
We investigate the constant-roll inflation with non-minimally kinetic coupling to the Einstein tensor. 
With the slow-roll parameter $\eta_\phi = -\ddot{\phi}/(H\dot{\phi})$ being a constant, 
we calculate the power spectra for scalar and tensor perturbations, 
and derive the expressions for the scalar spectral tilt $n_s$,
the tensor spectral tilt $n_T$, and the tensor-to-scalar ratio $r$.
We find that the expressions for $n_s$ are different with different ordering of taking the derivative of the scalar power spectrum with respect to the scale $k$ and the horizon crossing condition $c_sk=aH$ in the constant-roll inflation,
the consistency relation $r=-8n_T$ does not hold if $|\eta_\phi|$ is not small,
and the duality of the tensor-to-scalar ratio between the slow-roll inflation and ultra-slow-roll inflation does not exist in inflationary models with non-minimally derivative coupling.
The result offers a fresh perspective on the understanding of the inflationary models with non-minimally derivative coupling and is helpful for the production of scalar induced gravitational waves in the framework of ultra-slow-roll inflation with non-minimally derivative coupling.
\end{abstract}

\maketitle
\section{Introduction}
Inflation, characterized by an exponential expansion of the Universe during its early stages, was proposed to address issues in standard Big Bang cosmology, including the horizon, flatness, and monopole problems \cite{Guth:1980zm,Linde:1981mu,Albrecht:1982wi,Starobinsky:1980te,Guth:1982ec}. 
This expansion is typically driven by a scalar field known as the inflaton, where the simplest model is the canonical inflaton field with a flat potential.  
Utilizing the Higgs field, the only detected scalar field in the Standard Model of particle physics, as the inflaton leads to a conflict with observations of the cosmic microwave background (CMB) radiation due to the large prediction of the tensor-to-scalar ratio $r$. 
To overcome it, inflation models with the field non-minimally coupled to gravity have been proposed  \cite{Bezrukov:2007ep,Germani:2010gm, Yi:2018gse}.  
One notable model is the new Higgs inflation model with the kinetic term non-minimally derivative coupling to the Einstein tensor,  represented by  $G^{\mu\nu}\partial_\mu\phi\partial_\nu\phi/M^2$ \cite{Germani:2010gm,Germani:2014hqa}.  
Under the general relativity (GR) limit $F=H^2/M^2 =0$, this model becomes the canonical situation conflicting with the observations, 
while under the high friction limit $F\gg1$, the tensor-to-scalar ratio is $1/3$ of that in the canonical Higgs inflation, aligning with the observational data \cite{Germani:2014hqa,Yang:2015pga}. 
Furthermore, the effective self-coupling $\lambda$ of the Higgs boson in this model can be lowered to the order of unity without introducing a new degree of freedom  \cite{Germani:2010gm,Germani:2014hqa}.  For more works about the inflation model with  non-minimally derivative coupling, see  Refs. \cite{Germani:2010ux,Minamitsuji:2013ura,Sadjadi:2012zp,Zhu:2015lry,Yang:2015zgh,Gong:2017kim,Fu:2019ttf, Granda:2019wip,Oliveros:2019xef,Gialamas:2020vto, Heydari:2021gea,Goodarzi:2022wli,Oikonomou:2024tjf,Gialamas:2024jeb,Hell:2024xbv}. 

Apart from the slow-roll inflation scenario, the constant-roll inflation scenario, where one slow-roll parameter remains constant and need not be small, can also yield predictions consistent with observational data \cite{Martin:2012pe,Motohashi:2014ppa,Yi:2017mxs}. 
The characteristics of primordial perturbations in constant-roll inflation can differ significantly from those in slow-roll inflation. For instance, in the canonical constant-roll inflation with a constant $\eta_\phi = -\ddot{\phi}/(H\dot{\phi})$ and $\eta_\phi>3/2$, scalar perturbations outside the horizon continue to evolve  \cite{Motohashi:2014ppa}. Additionally, the power spectrum of scalar perturbations can exhibit sharp enhancements at small scales in ultra-slow-roll inflation with $\eta_\phi \approx 3$, potentially leading to the formation of primordial black holes and inducing secondary gravitational waves \cite{Germani:2017bcs,Gong:2017qlj,Ezquiaga:2017fvi, Ballesteros:2017fsr,Drees:2019xpp,Lin:2020goi, Lin:2021vwc, Yi:2020cut,Yi:2020kmq,Yi:2021lxc,Yi:2022anu,Zhang:2021rqs,Wang:2024euw}. 
Despite the differences in the evolution of perturbations between ultra-slow-roll inflation  $\eta_\phi =3-\alpha$  and slow-roll inflation  $\eta_\phi =\alpha$  in the canonical inflation model \cite{Gao:2019sbz}, if we neglect the contribution from $\epsilon_H$ to the scalar spectral tilt $n_s$, the predictions for the scalar spectral tilt $n_s$ and tensor-to-scalar ratio $r$ are the same \cite{Tzirakis:2007bf,Morse:2018kda}. This indicates the existence of a duality in constant-roll inflation models with large and small $\eta_\phi$. 
For more about the constant-roll inflation, please refer to Refs. \cite{Motohashi:2017vdc, Oikonomou:2017bjx, Odintsov:2017qpp,Nojiri:2017qvx, Dimopoulos:2017ged,Ito:2017bnn,Karam:2017rpw,Fei:2017fub,Cicciarella:2017nls,Anguelova:2017djf,Gao:2018tdb,Gao:2018cpp,Mohammadi:2018wfk, Pattison:2018bct,Lin:2019fcz,Fei:2020jab,Antoniadis:2020dfq,Guerrero:2020lng,Shokri:2021iqp,Fei:2022gfp,Stojanovic:2023dni,Huang:2024xqk}. 

In this paper, we research the constant-roll inflation model with non-minimally derivative coupling.   
We calculate the power spectrum of scalar and tensor fluctuations using the Bessel function method and derive expressions for the scalar spectral tilt $n_s$, the tensor spectral tilt $n_T$,  and the tensor-to-scalar ratio $r$. Additionally, we explore the dual between the ultra-slow-roll inflation and slow-roll inflation. 
The organization of this paper is as follows: Section \ref{sec:bg} presents the background evolution and defines the slow-roll parameters. Section \ref{sec:result} calculates the power spectrum for scalar and tensor perturbations using the Bessel function method within the constant-roll condition.  Section \ref{sec:dual} explores  the dual behavior between ultra-slow-roll inflation and slow-roll inflation. Finally, section \ref{sec:con} draws the conclusions.

\section{Inflation with non-minimally derivative coupling}\label{sec:bg}
In this section, we derive the background evolution, define the slow-roll parameters, and  discuss the general formulae for the the scalar and tensor spectral for the inflationary models with non-minimally derivative coupling. 
\subsection{The background}
The action for the inflationary model with the non-minimal derivative coupling is
\begin{eqnarray} \label{action}
S=\frac{1}{2}\int\text{d}^{4} x \sqrt{-g}\bigg[M_{p l}^{2} R-g^{\mu \nu} \partial_{\mu} \phi \partial_{\nu} \phi \nonumber\\
+\frac{1}{M^{2}} G^{\mu \nu} \partial_{\mu} \phi \partial_{\nu} \phi-2 V(\phi)\biggr],
\end{eqnarray}
where $M_{pl} =1/\sqrt{8\pi G}$ is the reduced planck mass,  $R$ is the Ricci scalar, $\phi$ is the scalar field, $G_{\mu\nu}$ is the Einstein tensor, $V(\phi)$ is the potential, and $M^2$ is the coupling parameter with the dimension of energy. 
For the homogeneous and isotropic background with the flat  Friedmann–Robertson–Walker (FRW) metric,  the background equations are \cite{Yang:2015pga}
\begin{gather} \label{bkeq1}
H^{2}=\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3 M_{pl}^{2}}\left[\frac{\dot{\phi}^{2}}{2}(1+9 F)+V(\phi) \right],\\
\label{bkeq2}
\frac{{\rm d}}{{\rm d} t}\left[a^{3} \dot{\phi}\left(1+3 F)\right)\right]=-a^{3} \frac{{\rm d} V}{{\rm d} \phi},\\
\label{Ray1}
\dot{H}-\frac{\dot{\phi} \ddot{\phi}}{M^{2} M_{p l}^{2}} H=\frac{1}{2} \frac{\dot{\phi}^{2}}{M^{2} M_{p l}^{2}} \dot{H}-\frac{3}{2} \frac{\dot{\phi}^{2}}{M^{2} M_{p l}^{2}} H^{2}-\frac{1}{2} \frac{\dot{\phi}^{2}}{M_{p l}^{2}},
\end{gather}
where $F=H^2/M^2$ and an over-dot means the derivative with respect to the cosmic time $t$.  

In this paper, we define the slow-roll parameters as 
\begin{equation}\label{srp}
 \epsilon_{H}=-\frac{\dot{H}}{H^2},\quad \eta_{H}=-\frac{\ddot{H}}{2H\dot{H}}, \quad \eta_{\phi}=-\frac{\ddot{\phi}}{H\dot{\phi}}.
 \end{equation}
In the canonical inflation models, the slow-roll parameters satisfy $\eta_H =  \eta_\phi$, a condition not met in the inflation models with non-minimal derivative coupling. By using the slow-roll parameters \eqref{srp}, the background equation \eqref{Ray1} can be rewritten as
\begin{equation}\label{Ray2}
 - \epsilon_{H}+2\eta_{\phi}Q=-\epsilon_{H}Q-3Q-Q/F,
\end{equation}
where the definition of $Q$ is
\begin{equation}
\label{qdef1}
Q=\frac{\dot{\phi}^{2}}{2 M^{2} M_{pl}^{2}}.
\end{equation}
Solving the background equation \eqref{Ray2}, we get  
\begin{equation}
\label{A} 
Q=\frac{F \epsilon_{H}}{b+F \epsilon_{H}},
\end{equation}
where $b=1+3 F+2 \eta _{\phi}F$.

For later use, we also give the following relation,
\begin{equation}
\label{dotepsh}
\dot{\epsilon}_H=2H(\epsilon_H-\eta_H)\epsilon_H,    
\end{equation}
\begin{equation}
\label{dotfeq}   
\dot{F}=-2HF\epsilon_H,
\end{equation}
and
\begin{equation}\label{dah}
\frac{\text{d}}{\text{d} \tau}\left(\frac{1}{a H}\right)=-1+\epsilon_{H},
\end{equation}
where the conformal time $\tau$ is related to the cosmic time $t$ by $\text{d} t=a \text{d} \tau$.

\subsection{The perturbations}
In the uniform field gauge $\delta \phi(x,t) =0$,  the action for the primordial  scalar perturbation  $\zeta$ is   \cite{Germani:2010ux,Yang:2015pga, Germani:2011ua} 
\begin{equation}\label{action3}
    S_{\zeta^{2}}=\int \text{d} t \text{d}^{3} x M_{pl}^{2} a^{3}\left\{\frac{\Sigma}{\bar{H}^{2}} \dot{\zeta}^{2}-\frac{\theta_{s}}{a^{2}}\left(\partial_{i} \zeta\right)^{2}\right\},
\end{equation}
where 
\begin{eqnarray}\label{1}
 \bar{H}=\frac{H(1-3 Q )}{1-Q }, \quad\Sigma=Q M^{2}\left[1+\frac{3F(1+3 Q)}{(1-Q)}\right],
\end{eqnarray}
and 
\begin{equation}
    \theta_{s}=\frac{1}{a} \frac{\text{d}}{\text{d} t}\left[\frac{a}{\bar{H}}(1-Q)\right]-1-Q.
\end{equation}
By the variation principle, we can derive the equation of motion  governing  the primordial perturbation $\zeta$; and in the Fourier space, we get the   Mukhanov-Sasaki equation 
\begin{equation}\label{ms}
v_{k}^{\prime \prime}+\left(c_{s}^{2} k^{2}-\frac{z^{\prime \prime}}{z}\right) v_{k}=0,
\end{equation}
where the prime denotes the derivative with respect to the conformal time $\tau$,
the effective sound speed is $c_{s}^{2}=\bar{H}^{2} \theta_{s} / \Sigma$, the canonically normalized field is $v=z \zeta$, and 
\begin{equation}\label{z}
    z=a M_{p l} \frac{\sqrt{2 \Sigma}}{\bar{H}}. 
\end{equation}
By using the slow-roll parameters \eqref{srp} and equation \eqref{Ray2},  we have 
\begin{eqnarray}\label{cs}
c_{s}^{2}=&&1-\frac{2 \epsilon_{H}F(1+7 F+6 F \eta_{\phi})}{(1+3 F)^{2}+2 F(1+3 F)\eta_{\phi}+12 \epsilon_{H}F^{2}} \nonumber\\
&&-\frac{8\epsilon_{H}^{2}F^3}{b[b+3F(b+4F\epsilon_{H})]},
\end{eqnarray}
and 
 \begin{align}   
 \label{ztoz}
   \frac{z^{\prime \prime}}{z}
=&a^2H^{2}\bigg[2+3 \eta_{H}-4\epsilon_{H}-6 \eta_{\phi}-\eta_{\phi}(2 \epsilon_{H}-2 \eta_{H}+\eta_{\phi})\nonumber\\
&+\frac{\eta_{H}-\eta_{\phi}}{F}\biggr]+\frac{a^{2} H^{2}b\epsilon_{H}(3-2 \eta_{\phi}+3 \epsilon_{H}-2 \eta_{H})}{b+3 F(b+4 F \epsilon_{H})}\nonumber\\
&+\frac{8 a^{2} H^{2} F \epsilon_{H} m(b+F \epsilon_{H})^{2}}{b^{2}\left(b-2 F \epsilon_{H}\right) \left(b+4 F \epsilon_{H}\right)}\nonumber\\
&+\frac{48 a^{2} H^{2} F^{2} \epsilon_{H}^{2} \eta_{\phi}^{2}(b+F \epsilon_{H})}{b(b-2 F \epsilon_{H})^{2}}\nonumber\\
&+\frac{48 a^{2} H^{2} F^{3} \eta_{\phi}^{2}\epsilon_{H}^{3}(b+F\epsilon_{H})}{b^{2}(b+4 F \epsilon_{H})^2}\nonumber\\&-\frac{a^{2} H^{2} \epsilon_{H}^{2}[4 F\eta_{\phi} (b+F \epsilon_{H})+b(b+4F\epsilon_{H})]^{2}}{(b+4 F \epsilon_{H})^{2}[b+3 F(b+4 F \epsilon_{H})]^{2}}\nonumber\\
 &-\frac{4 a^{2} H^{2} F \epsilon_{H}(b+F \epsilon_{H})}{b+3 F(b+4 F \epsilon_{H})}\bigg[\frac{2 \eta_{\phi} \epsilon_{H}}{b-2 F \epsilon_{H}}
 +\frac{\eta_{H}-\eta_{\phi}}{F(b+4 F \epsilon_{H})}
 \nonumber\\
 &+\frac{3(\eta_{H}-\epsilon_{H}-2 \eta_{\phi})-2 \eta_{\phi}(2\epsilon_{H}-\eta_{\phi}-\eta_{H})}{b+4 F \epsilon_{H}}\nonumber\\
 &- \frac{4 F\epsilon_{H} \eta_{\phi}^{2}(b-2 F\epsilon_{H})}{b(b+4 F\epsilon_{H})^{2}}\nonumber\\
 &+\frac{8F\eta_{\phi}^{2} \epsilon_{H}(b+F\epsilon_{H})}{b(b-2 F\epsilon_{H})(b+4 F\epsilon_{H})}\biggr],
\end{align}
where
\begin{align}
m&=2\eta_{\phi}(\eta_{\phi}+\epsilon_{H})+4F \eta_{\phi}^{2} (\eta_{H} +\eta_{\phi})+(1+3 F)\times
\nonumber\\
&
\bigg[3(\eta_{H}-\epsilon_{H}-2 \eta_{\phi})
-2 \eta_{\phi}(2 \epsilon_{H}+\eta_{\phi}-2 \eta_{H})+\frac{\eta_{H}-\eta_{\phi}}{F}\biggr].
\end{align}
The definition of the power spectrum of the scalar perturbations is
\begin{equation}  \label{def:spectrumS}
\mathcal{P}_{\zeta}  =\frac{k^{3}}{2 \pi^{2}}\left|\zeta_{k}\right|^{2},
\end{equation} 
and the scalar spectral tilt $n_s$ is defined by \cite{Kinney:2005vj}
\begin{equation}\label{def:ns}
  n_s-1=\left.\frac{{\rm d}\ln \mathcal P_{\zeta}}{{\rm d} \ln k}\right|_{a H =\text{const}}.
\end{equation}

For the tensor perturbation,  the  quadratic action  is \cite{Yang:2015pga, Germani:2011ua}
\begin{equation}\label{actiont}
    S=\int \text{d}^{3} x \text{d} t \frac{M_{p l}^{2}a^{3}}{8} \left[(1-Q) \dot{\gamma}_{i j}^{2}-\frac{1}{a^{2}}(1+Q)(\partial_{l} \gamma_{i j})^{2}\right].
\end{equation}
The   Mukhanov-Sasaki equation   for tensor perturbation is
\begin{equation}\label{ms:gw}
    u_{k}^{s \prime \prime}+\left(c_{t}^{2} k^{2}-\frac{z_t''}{z_t}\right) u_{k}^{s}=0.
\end{equation}
Here,   $u^{s} = z_t \gamma^s$, the label $s =\{ +, \times\}$, denotes the polarization of the tensor perturbations,  and 
\begin{equation}
   \gamma_{i j}=\sum_{s=+, \times} \text{e}_{i j}^{s} \gamma^{s},
\end{equation}
where $\text{e}_{i j}^{s}$ is the polarizaiton tensor and  satisfies  $\sum_{i} \text{e}_{i i}^{s}=0$, $\sum_{i, j} \text{e}_{i j}^{s}\text{e}_{i j}^{s^{\prime}}=2 \delta_{s s^{\prime}}$.   The sound speed  in equation \eqref{ms:gw} is 
\begin{equation}
    c_{t}^{2}=\frac{1+Q}{1-Q},
\end{equation}
and the normalized parameter is 
\begin{equation}
    z_{t}=\frac{\sqrt{2}}{2} a M_{p l} \sqrt{1-Q}.
\end{equation}
By using the slow-roll parameters \eqref{srp},  equation \eqref{Ray2}, and the relation \eqref{A}, we obtain
\begin{equation}\label{ct}
    c_{t}^{2}=  1+\frac{2F\epsilon_{H}}{1+(3+2\eta_{\phi})F},
\end{equation}
and
\begin{align} 
\label{zt}
     \frac{z_{t}^{\prime \prime}}{z_{t}}=
     &a^{2} H^{2}\left[2-\epsilon_{H}-\frac{F\epsilon_{H}}{1+(3+2\eta_{\phi})F}\bigg(\frac{\eta_{H}-\eta_{\phi}}{F}\right.\nonumber\\
     &+3(\eta_{H}-2\eta_{\phi}-\epsilon_{H})
     -2\eta_{\phi}(\epsilon_{H}-\eta_{H})
     \bigg)\nonumber\\
     &-\frac{F^2\epsilon_{H}^2\eta_{\phi}^2}{[1+(3+2\eta_{\phi})F]^2}\biggr].
\end{align}
The definition of the power spectrum for the tensor perturbations is
\begin{equation}\label{def:Pt}
 \mathcal{P}_{T}   =\frac{k^{3}}{2 \pi^{2}}\left|\gamma_{i j}\right|^{2}=\frac{k^{3}}{\pi^{2}} \sum_{s=+, x}\left|\frac{u_{k}^{s}}{z_{t}}\right|^{2}.
\end{equation}


\section{The Constant-roll inflation} \label{sec:result}  
In addition to the slow-roll inflation scenario, where all the slow-roll parameters are small, the constant-roll inflation scenario, where one of the slow-roll parameters is held constant and need not be small, can also produce predictions consistent with observational data from  CMB.  In this section, we  study the constant-roll scenario where the slow-roll parameter $\eta_\phi$ is a constant.   

Combining equations \eqref{qdef1} and \eqref{A} and calculating $\dot{Q}$, we get 
\begin{equation}\label{dphi}
\begin{split}
   \dot{ \eta_{\phi}}=&3H( \eta_{\phi}- \eta_{H}+\epsilon_{H})+ \eta_{\phi}H(3\epsilon_{H}-2 \eta_{H}+2 \eta_{\phi})\\
   &-\frac{( \eta_{H}- \eta_{\phi})H}{F}.
\end{split}
\end{equation}
By using the condition that the slow-roll parameter $\eta_\phi$ is a constant,  we obtain
\begin{equation} \label{etahrel1}
\eta_H=\eta_\phi+\frac{3F\epsilon_H(\eta_\phi+1)}{1+(3+2\eta_\phi)F},   
\end{equation}
and relation \eqref{dotepsh} becomes, 
 \begin{equation}\label{depsilon}
    \dot{\epsilon}_H=2H\epsilon_H\left(\epsilon_H-\eta_\phi-\frac{3F\epsilon_H(\eta_\phi+1)}{1+(3+2\eta_\phi)F}\right).  
\end{equation}   
Because $\eta_\phi$ is a constant and it may not be small,
so $\dot{\epsilon}_H$ can still be the first order in perturbation.
 
In the GR limit, $F=0$, we recover the standard result $\eta_H=\eta_\phi$.
In the high friction limit, $F\gg 1$,
\begin{equation}
\label{etahrel2}   
\eta_H=\eta_\phi+\frac{3\epsilon_H(\eta_\phi+1)}{3+2\eta_\phi}.
\end{equation}
 Because the slow-roll parameter $\eta_\phi$ may not be small in the constant-roll inflation scenario, 
 the slow-roll parameter $\eta_H$ may not be small in general either.

For the constant-roll inflation with constant $\eta_\phi$, from equation \eqref{dah} and to the first order approximation of $\epsilon_{H}$, we get the relation of the conformal Hubble parameter $aH$ and the conformal time $\tau$ \cite{Yi:2017mxs},
\begin{equation}
\label{ah}
\frac{1}{aH}\approx\left(-1+\frac{\epsilon_{H}}{1+2\eta_{\phi}}\right)\tau.
\end{equation}
\subsection{Scalar Perturbations} 
To the first order of  $\epsilon_{H}$,  the relation \eqref{ztoz} becomes 
\begin{equation}
\begin{split}
\label{czz}
    \frac{z''}{z} \approx & a^{2} H^{2}\bigg[2-3 \eta_{\phi}+\eta_{\phi}^{2}+\frac{2-3 F}{1+3 F} \epsilon_{H}\\
    & -\eta_{\phi} \epsilon_{H}\left(\frac{3(1-F)}{1+3 F}  +\frac{4 F (1+6 F)(3 -4 \eta_{\phi})}{(1+3 F)[1+(3+2 \eta_{\phi}) F]}\right)\biggr].
 \end{split}     
 \end{equation} 
To solve the Mukhanov-Sasaki equation \eqref{ms},  
we write  
\begin{equation}\label{nu}
\frac{z''}{z}=\frac{\nu^2-1/4}{\tau^2}.
\end{equation} 
To the first order of  $\epsilon_{H}$,  
combining equations \eqref{ah}  and \eqref{czz},  we get
 \begin{eqnarray} 
 \label{nu1}
 \nu&=&\left|\frac{3}{2}-\eta_{\phi} \right|+\frac{\bar\nu \epsilon_{H}}{\left|3/2-\eta_{\phi}\right|},
\end{eqnarray}
where    
\begin{equation}
\begin{split}\label{nubar}
\bar\nu=&\frac{2-3 \eta_{\phi}+\eta_{\phi}^{2} }{1+2 \eta_{\phi}}+\frac{2-3 F}{2(1+3 F)} \\
&-\eta_{\phi}\left[\frac{3(1-F)}{2(1+3 F)} +\frac{2 F(1+6 F) (3-4 \eta_{\phi}) }{(1+3 F)[1+(3+2 \eta_{\phi}) F]}\right].
\end{split}
\end{equation}
Assuming that both $\nu$ and $c_s$ are constants, 
then the Mukhanov-Sasaki equation \eqref{ms} becomes the standard Bessel equation, 
and the solution can be expressed as the Hankel function of order $\nu$ \cite{Stewart:1993bc},
\begin{equation}\label{Hankel:func}
     v_{k}(\tau)=\sqrt{-\tau}\left[c_{1} H_{\nu}^{(1)}(-c_s k \tau)+c_{2} H_{\nu}^{(2)}(-c_s k \tau)\right] . 
\end{equation}
Choosing the  Bunch-Davies vacuum, 
at small scales, $-c_s k \tau \gg1$, we have the initial condition
\begin{equation}\label{vk0}
v_{k} \rightarrow \frac{1}{\sqrt{2 c_{s} k}} \text{e}^{\text{-i} c_{s} k \tau}. 
\end{equation}
Using the inital condition \eqref{vk0} and the asymptotic behavior of the Hankel function,  
we obtain the mode function outside the horizon
\begin{equation}\label{sol:vk}
    v_{k}=\text{e}^{\text{-i}(\nu-1 / 2) \frac{\pi}{2}} 2^{\nu-\frac{3}{2}} \frac{\Gamma(\nu)}{\Gamma(3 / 2)} \frac{1}{\sqrt{2 c_{s} k}}\left(-c_{s} k \tau\right)^{\frac{1}{2}-\nu}.
\end{equation}
Combining equation \eqref{z} and the solution \eqref{sol:vk}, from the definition of the  power spectrum for the scalar perturbation  \eqref{def:spectrumS}, we obtain  
\begin{eqnarray}
  \label{spectrumS}
\mathcal{P}_{\zeta}  
 \approx 2^{2 \nu-3}\left[\frac{\Gamma(\nu)}{\Gamma(3 / 2)}\right]^{2} \frac{H^{2}}{8 \pi^{2}M_{p l}^{2} c_{s} \theta_{s} } \nonumber\\
 \times\left(1-\frac{\epsilon_H}{1+2\eta_\phi}\right)^{2\nu-1}\left(\frac{c_{s}k}{aH}\right)^{3-2 \nu}.
\end{eqnarray} 
Substituting  equation \eqref{spectrumS} into the definition of the scalar spectral tilt  $n_s$, equation \eqref{def:ns}, we obtain 
\begin{equation}
\label{NS}
  n_s-1=  3-2\nu.
\end{equation}
If $|\eta_\phi|\ll 1$, to the first order of perturbation, we get \cite{Tsujikawa:2012mk,Yang:2015pga}
\begin{equation}
\label{srnsres}
n_s-1\approx 2\eta_\phi-\frac{2(2+3F)}{1+3F}\epsilon_H.   
\end{equation}

In the GR limit,  we  have
\begin{equation}
\nu=\left|\frac{3}{2}-\eta_{\phi}\right|+\frac{(6-5\eta_{\phi}-4\eta^2_{\phi})\epsilon_{H}}{\left|3-2\eta_{\phi}\right|(1+2\eta_{\phi})},
\end{equation}
and 
\begin{equation}
\label{grns1}
n_s-1=3-\left|3-2\eta_{\phi}\right|-\frac{2(6-5\eta_{\phi}-4\eta^2_{\phi})\epsilon_{H}}{\left|3-2\eta_{\phi}\right|(1+2\eta_{\phi})}.
\end{equation}

In the high friction limit,  we  have
\begin{equation}
    \nu=\left|\frac{3}{2}-\eta_{\phi}\right|+\frac{3(3-4\eta_{\phi})(1-3\eta_{\phi }-6\eta^2_{\phi})\epsilon_{H}}{\left|3-2\eta_{\phi}\right|(1+2\eta_{\phi})(3+2\eta_{\phi})},
\end{equation}
\begin{equation}
\begin{split}
n_s-1=&3-\left|3-2\eta_{\phi}\right|\\
&-\frac{6(3-4\eta_{\phi})(1-3\eta_{\phi }-6\eta^2_{\phi})\epsilon_{H}}{\left|3-2\eta_{\phi}\right|(1+2\eta_{\phi})(3+2\eta_{\phi})}.
\end{split}
\end{equation}

As discussed above, $\dot{\epsilon}_H$ is still the first order, 
so we will get a different expression for the scalar spectral tilt if we take the derivative of the amplitude of the scalar power spectrum at the horizon crossing $c_sk=aH$ with respect to the scale $k$. 
To show this point, we give the derivation below.

For the constant-roll inflation with large $\eta_\phi$, the scalar perturbations may continue evolving outside the horizon \cite{Motohashi:2014ppa}. If scalar perturbations are constant after the horizon crossing, 
$aH\geq c_s k$,  
to the first order of $\epsilon_{H}$, 
the  amplitude of the power spectrum for scalar perturbations outside the horizon is  
\begin{equation}
\begin{aligned}
 \label{Ps1}
\mathcal{A}_\zeta &= \mathcal{P}_{\zeta} |_{c_{s} k=a H} \\
&\approx  2^{|3-2 \eta_{\phi }|-3}\left[\frac{\Gamma(|3/2-\eta_\phi|)}{\Gamma(3 / 2)}\right]^{2} \frac{H^{2}}{8 \pi^{2}  M_{p l}^{2}\epsilon_{H}} \\
&\times \frac{1+3F+2F\eta_\phi}{1+3F}
 (1+D \epsilon_H), 
\end{aligned}
\end{equation}
where 
\begin{eqnarray}
\label{cdef1}
 D=&  &\frac{4\bar\nu}{|3-2 \eta_{\phi } |}\left(\ln2+\frac{\Gamma'(|3/2-\eta_\phi|)}{\Gamma(|3/2-\eta_\phi|)}\right)  
  -\frac{|3-2 \eta_{\phi } |-1}{1+2\eta_{\phi }}\nonumber\\
  &&\frac{18 F^{2} \eta_{\phi}}{(1+3 F)[1+(3+2\eta_{\phi})F]}.
\end{eqnarray}
Besides the definition \eqref{def:spectrumS}, 
if scalar perturbations are constants after the horizon crossing, 
the  scalar spectral tiltcan also be derived by using the  the  amplitude of the power spectrum for the scalar perturbation,
\begin{equation}\label{def:spectrumS2}
    n_s-1=\left.\frac{\text{d} \ln \mathcal{A}_{\zeta}}{\text{d} \ln k}\right|_{c_{s} k=a H}.
\end{equation}
To the first order of approximation of $\epsilon_H$, 
\begin{equation}
\label{dlnkeq}
\begin{split}
\frac{\text{d}\ln k}{H\text{d}t}&=(1-\epsilon_{H})-H^{-1}\frac{\dot{c_{s}}}{c_{s}}\\
    &\approx 1-\left(\frac{2 F\eta_{\phi}(1+7F+6F\eta_{\phi})  }{(1+3 F)[1+(3 +2  \eta_{\phi})F]}+1\right)\epsilon_{H}. \\
\end{split}
\end{equation}
Substituting  equations \eqref{Ps1} and \eqref{dlnkeq} into equation \eqref{def:spectrumS2},  
to the first order of $\epsilon_H$,  
we get
\begin{equation}
\begin{aligned}
\label{lnPb}
n_s&-1 \approx2 \eta_{\phi}-\frac{2(2+3F)\epsilon_{H} }{1+3 F}\\
&+2\eta_{\phi}\epsilon_{H}  \Bigg[ \frac{|3-2\eta_{\phi}|-1}{1+2\eta_{\phi}} \\
&-\frac{4\Bar{\nu} }{\left|3-2\eta_{\phi}\right|}\left(\ln2+\frac{\Gamma'(|3/2-\eta_\phi|)}{\Gamma(|3/2-\eta_\phi|)}\right)\\
&+\frac{(1+3F)(1+4F)+2F\eta_{\phi }(2+F+6F\eta_{\phi })}{(1+3 F)[1+(3+2  \eta_{\phi})F]}\biggr].    
\end{aligned}
\end{equation}
It is obvious this result is different from equation \eqref{NS} if $\eta_\phi$ is not small. 
If $|\eta_\phi|\ll 1$, then the result \eqref{lnPb} reduces to the result \eqref{srnsres}.
In the following discussion,
we use the familiar result \eqref{NS}.

\subsection{Tensor Perturbation}
For tensor perturbations, to solve equation \eqref{ms:gw}, 
we use the same method as that solving the scalar perturbation.  
To the first order of  $\epsilon_{H}$, we have
\begin{equation}
\label{zt2}
\begin{split}
\frac{z_{t}^{\prime \prime}}{z_{t}} &\approx  a^{2} H^{2}\left[2-\epsilon_{H}+\frac{F\epsilon_{H}\eta_{\phi}(3-2\eta_{\phi})}{1+(3+2\eta_{\phi})F}\right]\\
&= \frac{\mu^2-1/4}{\tau^2},
\end{split}
\end{equation}
where
\begin{equation}\label{mu2}
    \mu \approx \frac{3}{2}+\bar\mu \epsilon_{H},
\end{equation}
and
\begin{equation}
  \bar\mu=  \frac{(3-2 \eta_{\phi})[1+3F+F\eta_{\phi}(3+2 \eta_{\phi})]}{3(1+2 \eta_{\phi})[1+(3+2  \eta_{\phi})F]} .
\end{equation}

Like scalar perturbations, assuming that both $c_t$ and $\mu$ are constants,   
to the first order of $\epsilon_{H}$, 
we obtain the power spectrum for the tensor  perturbation,
\begin{equation}\label{Pt}
\begin{aligned}
  \mathcal{P}_{T}  & 
   \approx \frac{2^{2\mu}}{M_{pl}^{2}}\left(\frac{H}{2\pi}\right)^{2}\left[\frac{\Gamma(\mu)}{\Gamma(3 / 2)}\right]^{2}\left(1-\frac{\epsilon_H}{1+2\eta_{\phi}}\right)^{2\mu-1}\\ &\left(\frac{1+3F+2F\eta_{\phi}-2F\epsilon_{H}}{1+(3+2\eta_{\phi})F}\right)\left(\frac{c_tk}{aH}\right)^{(3-2\mu)}.
\end{aligned}
\end{equation}
Therefore, the tensor spectral tilt is
\begin{equation}
\label{NT}
\begin{split}
n_T&=\left.\frac{\text{d} \ln \mathcal P_{T}}{\text{d} \ln k}\right|_{a H =\text{const}}\\
&=-\frac{2(3-2 \eta_{\phi})[1+3F+F\eta_{\phi}(3+2 \eta_{\phi})]}{3(1+2 \eta_{\phi})[1+(3 +2  \eta_{\phi})F]}\epsilon_H.
\end{split}
\end{equation} 
If $|\eta_\phi|\ll 1$, then we get $n_T\approx-2\epsilon_H$.

In the GR limit, we get
\begin{equation}
\label{NT}
n_T=-\frac{2(3-2 \eta_{\phi})}{3(1+2 \eta_{\phi})}\epsilon_H.
\end{equation} 
In the high friction limit, we get
\begin{equation}
\label{NT}
n_T=-\frac{2(3-2 \eta_{\phi})[3+\eta_{\phi}(3+2 \eta_{\phi})]}{3(1+2 \eta_{\phi})(3+2 \eta_{\phi})}\epsilon_H.
\end{equation} 

To the first order of $\epsilon_{H}$, 
the amplitude of the power spectrum for 
tensor perturbations becomes
\begin{eqnarray}\label{pt2}
\mathcal{P}_{T}\bigg|_{c_{t} k=a H}
\approx\frac{8}{M_{pl}^2}\left(\frac{H}{2\pi}\right)^{2}\bigg[1-(A-B)\epsilon_{H}\biggr],
\end{eqnarray}

where
\begin{equation}
\begin{split}
A&=\frac{2[3+C(3-2\eta_{\phi})]}{3(1+2\eta_{\phi})},\\
B&=-\frac{2F[C\eta_{\phi}(3-2\eta_{\phi})+3]}{3[1+(3+2\eta_{\phi})F]},
\end{split}
\end{equation}
and  the constant $C=-2+\gamma+\ln{2}\approx-0.73 $. 

Combining the scalar power spectrum \eqref{Ps1} and the tensor power spectrum \eqref{pt2},
we obtain the  tensor-to-scalar ratio 
\begin{equation} \label{r1}
\begin{aligned}
r &=\frac{\mathcal{P}_T|_{c_{t} k=a H}}{\mathcal{P}_\zeta|_{c_{s} k=a H}} \\
 &=\frac{2^{7-|3-2\eta_{\phi}|} [\Gamma(3/2)]^2 (1+3F) }{\left[\Gamma(|3/2-\eta_{\phi}|)\right]^2[1+(3+2\eta_{\phi})F]} \epsilon_H.  
\end{aligned}
\end{equation}
It is apparent that the consistency relation $r=-8n_T$ does not hold for the constant-roll inflation if $|\eta_\phi|$ is not small.
If $|\eta_\phi|\ll 1$, then 
\begin{equation}
\label{r4}
r\approx 16 \epsilon_H,
\end{equation}
the consistency relation still holds.

In the GR limit, we get 
\begin{equation}
\label{r2}
r=\frac{2^{7-|3-2\eta_{\phi}|}[\Gamma(3/2)]^2}{\left[\Gamma(|3/2-\eta_{\phi}|)\right]^2}\epsilon_H.
\end{equation}
In the high friction limit, we get
\begin{equation}
\label{r3}
r=\frac{3 [\Gamma(3/2)]^2 2^{7-|3-2\eta_{\phi}|} }{\left[\Gamma(|3/2-\eta_{\phi}|)\right]^2(3+2\eta_{\phi})} \epsilon_H.
\end{equation}

\section{The duality in constant-roll inflation}\label{sec:dual}

In the GR case, from equation \eqref{grns1}, 
we see that if we neglect the contribution from $\epsilon_H$,
then we have the same scalar spectral tilt $n_s-1\approx 2\alpha$
when we replace $\eta_\phi=\alpha$ by $\eta_\phi=3-\alpha$ with $|\alpha|<3/2$.
From equation \eqref{r2}, we see that the duality between $\eta_\phi=\alpha$ and $\eta_\phi=3-\alpha$ also exists for the tensor-to-scalar ratio $r$.
The duality is also referred to as the duality between constant slow-roll and ultra-slow-roll inflation   and explained as the two branches of the solution of the  inflation field from the same potential \cite{Morse:2018kda} 
although the behaviors of background and perturbations are different for the constant slow-roll and ultra-slow-roll inflationary models if $\epsilon_H$ is not negligible \cite{Gao:2019sbz}. 
In this section, we discuss this duality behavior, 
aiming to understand the connection between the slow-roll and ultra-slow-roll scenarios within the framework of inflationary models with non-minimally derivative coupling.

Combining equations \eqref{nu1} and \eqref{NS}, and neglecting the contribution from $\epsilon_H$, we get
\begin{equation}
\label{dualns1}
n_s-1= 3-|3-2\eta_\phi|.
\end{equation}
It is obvious that the duality between $\eta_\phi=\alpha$ and $\eta_\phi=3-\alpha$  holds, and
\begin{equation}
 n_s-1  \approx 2 \alpha,
\end{equation}
if $|\alpha|<3/2$.
Now we discuss whether the duality holds for the tensor-to-scalar ratio $r$.
For the constant slow-roll inflation with $\eta_\phi=\alpha$, from equation \eqref{r1}, we get
\begin{equation}
\label{r6}
r=\frac{2^{4+2\alpha} [\Gamma(3/2)]^2 (1+3F) }{\left[\Gamma(3/2-\alpha)\right]^2[1+(3+2\alpha)F]} \epsilon_H.
\end{equation}
For the ultra-slow-roll inflation with $\eta_\phi =3 -\alpha$,  the tensor-to-scalar ratio becomes
\begin{equation}
\label{r7}
r=\frac{2^{4+2\alpha} [\Gamma(3/2)]^2 (1+3F) }{\left[\Gamma(3/2-\alpha)\right]^2[1+(9-2\alpha)F]} \epsilon_H.
\end{equation}
Comparing equations \eqref{r6} and \eqref{r7}, we find that they are the same only if $F=0$,
i.e., the duality exists only in the GR case, and there is no such duality in inflationary models with non-minimally derivative coupling.

When the slow-roll parameter $\epsilon_H$ is extremely small and can be neglected, 
the scalar spectral tilts obtained from both the slow-roll inflation and the ultra-slow-roll inflation coincide. 
In the limit where $F=0$ (GR limit), the tensor-to-scalar ratio \eqref{r6} derived from slow-roll inflation matches equation \eqref{r7} obtained from ultra-slow-roll inflation, 
thus the duality recovers in the GR limit.
However, in more general case, the tensor-to-scalar ratios $r$  from slow-roll inflation and ultra-slow-roll inflation are different. 
Consequently, even for the scalar spectral tilt $n_s$ and the tensor-to-scalar ratio $r$, 
the duality behavior between slow-roll and ultra-slow-roll inflation models with non-minimally derivative coupling is not observed. 
Therefore, the duality of $n_s$ and $r$ is not a universal feature in the constant-roll inflation models. 

\section{conclusion} 
\label{sec:con}

We study the constant-roll inflation model with the kinetic term non-minimally derivative coupling to the Einstein tensor. 
For the constant-roll condition, we choose the slow-roll parameter $\eta_\phi  =-\ddot{\phi}/( H\dot{\phi})$ to be a constant. 
With this constant-roll condition, we calculate the power spectra for the scalar and tensor perturbations and subsequently derive the scalar spectral tilt $n_s$, 
tensor spectral tilt $n_T$, and tensor-to-scalar ratio $r$. 
Due to the effect of large $\eta_\phi$, 
$\dot{\epsilon}_H$ remains to be the first order,
the expressions for $n_s$ are different with different ordering of taking the derivative of the scalar power spectrum with respect to the scale $k$ and the horizon crossing condition $c_sk=aH$ in the constant-roll inflation,
and the consistency relation between the tensor-to-scalar ratio and tensor spectral tilt, $r=-8n_T$, does not hold.
Between the slow-roll inflation with $\eta_\phi =\alpha$ being small and the ultra-slow-roll inflation with $\eta_\phi = 3-\alpha$, 
although the duality for the scalar spectral tilt still holds if we neglect the contribution of $\epsilon_H$ to $n_s$,
but the results for the tensor-to-scalar ratio $r$ are different. 
Thus, the duality of $n_s$ and $r$ is not a universal feature in inflation with non-minimally derivative coupling.
 
The inflation model with non-minimally derivative coupling is important because the new Higgs inflation model saves the Higgs inflation from unitarity bound violations. 
This investigation into constant-roll inflation presents a different perspective from the typical slow-roll inflation scenario, thereby enhancing our understanding of inflation models with non-minimally derivative coupling.  Scalar-induced gravitational waves usually originate from inflation models featuring a transitional ultra-slow-roll phase. Therefore, this investigation into the ultra-slow-roll condition is also helpful for the issue of scalar-induced gravitational waves in the inflation model with non-minimally derivative coupling.

\begin{acknowledgments}
JL is supported by the Hainan Provincial Natural Science Foundation of China under Grant No.121MS033. YG  is partially supported by  the National Key Research and Development Program of China under Grant No. 2020YFC2201504. ZY is supported by the National Natural Science Foundation of China under Grant No. 12205015.
\end{acknowledgments}


\end{document}